OpenStudy (anonymous):

Suppose that the functions u and w are defined as follows: u(x)=x-1 w(x)=2x^2-1 Find the following: (w o u)(-5)=? (u o w)(-5)=?

OpenStudy (jamesj):

So first of all, what are the two composite function (w o u)(x) and (u o w)(x) For example, suppose f(x) = x^2 and g(x) = 3x - 1. Then (f o g)(x) = f(g(x)) = f(3x - 1) = (3x-1)^2